March 11-14

– Three stores provided ~20.75 hours of luminosity
– Main Injector repaired transductor problem
– Antiproton Source power supply replaced
– The Tevatron quenched during shot setup due to another kicker problem
– Tevatron repaired higher order power supply problem
– Instrumentation technician repaired Tevatron BPM Front End
– Store 8571 lost due to a quench
– Tevatron sector B0 cryo system compressor failed

* The integrated luminosity for the period from March 7-14 was 6.18 inverse picobarns. NuMI is off for target replacement.

Read the Current Accelerator Update
Read the Early Bird Report
View the Tevatron Luminosity Charts